OpenStudy (anonymous):

What is another way to write the expression 5a + 4a + 2? A (5a + 4a) + 2 B 5a + 4a + 2a C 2 + a + 4 + a + 5 D (2 + 5a) + (2 + 4a)

OpenStudy (anonymous):

A

Parth (parthkohli):

\[\textbf{HINT}\]\[a + (b + c) = a + (b + c) = a + b + c\]